# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v3

@@ Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
$w_{t+1}=w_{t}+\alpha(r_{\Delta}(s^{t},a^{t})+\gamma\bar{\phi}_{k}(s^{t+1}_{\Delta})^{\top}w_{t}-\bar{\phi}_{k}(s_{\Delta}^{t})^{\top}w_{t})\bar{\phi}_{k}(s_{\Delta}^{t})\$ (43)
@@
$\hat{\Phi}_{k}=\begin{bmatrix}\bar{\Phi}_{k}(1)\\ \bar{\Phi}_{k}(2)\\ \vdots\\ \bar{\Phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|)\end{bmatrix}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}\$ (44)
Here, the system’s stage rewards are given by $r=\begin{bmatrix}r(1)&\dots&r(S)\end{bmatrix}\in\mathbb{R}^{S}$ and $r_{k}=[r_{k}(1),\dots,r_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|]\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|}$, where $d$ is a low-dimensional embedding we wish to learn.
@@
[…]
@@
Hence, in the continuous state/action space setting, as $k\to n$ and $M\to\infty$, $V^{\tilde{\pi}_{k}}(s_{0})\to V^{\tilde{\pi}_{n}}(s_{0})=V^{\pi^{*}}(s_{0})$. Intuitively, as $k\to n$, the optimality gap diminishes following the arguments in Corollary G.7. As $M\to\infty$, the number of samples obtained allows for more effective learning of the spectral features.

### v4

@@ Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
$w_{t+1}=w_{t}+\alpha(r_{\Delta}(s^{t},a^{t})+\gamma\bar{\phi}_{k}(s^{t+1}_{\Delta})^{\top}w_{t}-\bar{\phi}_{k}(s_{\Delta}^{t})^{\top}w_{t})\bar{\phi}_{k}(s_{\Delta}^{t})\\$ (45)
@@
$\hat{\Phi}_{k}=\begin{bmatrix}\bar{\Phi}_{k}(1)\\ \bar{\Phi}_{k}(2)\\ \vdots\\ \bar{\Phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|)\end{bmatrix}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}\\$ (46)
Here, the system’s stage rewards are given by
$r=\begin{bmatrix}r(1)&\dots&r(S)\end{bmatrix}\in\mathbb{R}^{S}$
and
$r_{k}=[r_{k}(1),\dots,r_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|]\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|},$
where $d$ is a low-dimensional embedding we wish to learn.
@@
$\mathbb{E}\left[\left|\hat{Q}_{k}^{\pi}(s,a)-\hat{Q}_{k}^{\mathrm{LFA}}(s,a)\right|\right]\leq O\left(\underbrace{\log\left(\frac{2k}{\delta}\right)\frac{\|\bar{\phi}_{k}\|^{5}}{\sqrt{M}}}_{\emph{statistical error}}+\underbrace{2\epsilon_{P}\gamma\cdot\frac{\tilde{r}}{1-\gamma}\cdot\|\bar{\phi}_{k}\|}_{\emph{approximation error}}\right),$
@@
Hence, in the continuous state/action space setting, as $k\to n$ and $M\to\infty$, $V^{\tilde{\pi}_{k}}(s_{0})\to V^{\tilde{\pi}_{n}}(s_{0})=V^{\pi^{*}}(s_{0})$. Intuitively, as $k\to n$, the optimality gap diminishes following the arguments in Corollary˜G.7. As $M\to\infty$, the number of samples obtained allows for more effective learning of the spectral features.
Appendix K Towards Deterministic Algorithms: Sharing Randomness
In large distributed systems, the random sampling in our communication network may be a bottleneck for “decentralized execution”. In light of this, we have provided a practical derandomized heuristic where the agents can share some randomness by only sampling within pre-defined fixed blocks of size. In this section, we propose algorithms RANDOM-SUB-SAMPLE-MFQ and RANDOM-SUB-SAMPLE-MFQ+, which shares randomness between agents through various careful groupings and using shared randomness within each groups. By implementing these algorithms, we derive simulation results, and establish progress towards derandomizing inherently stochastic approximately-optimal policies.
Algorithm˜5 (Execution with weakly shared randomness). The local agents are heuristically divided into arbitrary groups of size $k$. For each group $h_{i}$, the k1k-1 local agents sampled are the same at each iteration. The global agent’s action is then the majority global agent action determined by each group of local agents. At each round, this requires $O(\left\lceil n^{2}/k\right\rceil(n-k))$ random numbers.
Algorithm 5 SUB-SAMPLE-MFQ: Execution with weakly shared randomness
0: A multi-agent system as described in Section˜2. A distribution $s_{0}$ on the initial global state $s_{0}=(s_{g},s_{{[n]}})$. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})$.
1: Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$.
2: Define groups $h_{1},\dots,h_{x}$ of agents where $x\coloneq\left\lceil\frac{n}{k}\right\rceil$ and $|h_{1}|=|h_{2}|=\dots=|h_{x-1}|=k$ and $|h_{x}|=n\mod k$.
3:for $t=0$ to $T^{\prime}-1$ do
4:  for $i\in[x]$ do
5:   Let $\Delta_{i}$ be a uniform sample of $n]\setminus g_{i}\choose k-1$}, and let $a_{g_{i}}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta_{i}}(t))]_{1:k}$.
6:  Let $a_{g}(t)=\mathrm{majority}\left(\left\{[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta_{i}}(t))]_{g}\right\}_{i\in[x]}\right)$.
7:  Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t))$
8:  Let $s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t)$
9:  $s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
10:  Get reward $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
Algorithm˜6 (Execution with strongly shared randomness). The local agents are randomly divided in to groups of size $k$. For each group, each agent uses the other agents in the group as the k1k-1 other sampled states. Similarly, the global agent’s action is then the majority global agent action determined by each group of local agents. Here, at each round, this requires $O(\left\lceil n^{2}/k\right\rceil)$ random numbers.
Algorithm 6 SUB-SAMPLE-MFQ: Execution with strongly shared randomness
0: A multi-agent system as described in Section˜2. A distribution $s_{0}$ on the initial global state $s_{0}=(s_{g},s_{{[n]}})$. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})$.
1: Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$.
2: Define groups $h_{1},\dots,h_{x}$ of agents where $x\coloneq\left\lceil\frac{n}{k}\right\rceil$ and $|h_{1}|=|h_{2}|=\dots=|h_{x-1}|=k$ and $|h_{x}|=n\mod k$.
3:for $t=0$ to $T^{\prime}-1$ do
4:  for $i\in[x-1]$ do
5:   Let $a_{h_{i}}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{i}}(t))]_{1:k}$.
6:   Let $a^{g}_{h_{i}}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{i}}(t))]_{g}$.
7:  Let $\Delta_{\mathrm{residual}}\coloneq{[n]\setminus h_{x}\choose k-(n-\left\lceil n/k\right\rceil k)}$.
8:  Let $a_{h_{x}}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{x}\cup\Delta_{\mathrm{residual}}}(t))]_{1:n-\left\lceil n/k\right\rceil k}$
9:  Let $a^{g}_{h_{x}}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{x}\cup\Delta_{\mathrm{residual}}}(t))]_{g}$
10:  Let $a_{g}(t)=\mathrm{majority}\left(\{a_{h_{i}}^{g}(t):i\in[x]\}\right)$.
11:  Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t)),s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
12:  Get reward $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
The probability of a bad event (the policy causing the $Q$-function to exceed the Lipschitz bound) scales with the $O(n^{k})$ independent random variables, which is polynomially large. Agnostically, some randomness is always needed to avoid periodic dynamics that may occur when the underlying Markov chain is not irreducible. In this case, an adversarial reward function could be designed such that the system does not minimize the optimality gap, thereby penalizing excessive determinism.
This ties into an open problem of interest, which has been recently explored [Larsen et al., 2024]. What is the minimum amount of randomness required for SUBSAMPLE-MFQ (or general mean-field or multi-agent RL algorithms) to perform well? Can we derive a theoretical result that demonstrates and balances the tradeoff between the amount of random bits and the performance of the subsampled policy when using the greedy action from the $k$-agent subsystem derived from $\hat{Q}_{k}^{*}$? We leave this problem as an exciting direction for future research.